A simple random sample of 50 items from a population with standard dev=7  resulted in a sample mean of 36.

If required, round your answers to two decimal places.

a. Provide a 90% confidence interval for the population mean.
to

b. Provide a 95% confidence interval for the population mean.
to

c. Provide a 99% confidence interval for the population mean.
Answer #1

Solution :

Given that,

\bar x = 36

\sigma = 7

n = 50

a ) At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

Margin of error = E = Z\alpha/2* (\sigma\sqrt/n)

= 1.645 * (7 / \sqrt 50  )

= 1.63

At 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

36 - 1.63< \mu < 36 + 1.63

34.37< \mu < 37.63

(34.37, 37.63 )

b ) At 95% confidence level the z is ,

\alpha  = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.960

Margin of error = E = Z\alpha/2* (\sigma\sqrt/n)

= 1.960 * (7 / \sqrt 50  )

= 1.94

At 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

36 - 1.94< \mu < 36 + 1.94

34.06< \mu < 37.94

(34.06, 37.94 )

c ) At 99% confidence level the z is ,

\alpha  = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 2.576

Margin of error = E = Z\alpha/2* (\sigma\sqrt/n)

= 2.576 * (7 / \sqrt 50  )

= 2.55

At 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

36 - 2.55< \mu < 36 + 2.55

33.46< \mu < 38.55

(33.45, 38.55 )